OpenStudy (anonymous):

What is the rate constant of a first-order reaction that takes 5.50 minutes for the reactant concentration to drop to half of its initial value?

OpenStudy (dumbcow):

im assuming this is a decay function 5.5 is the half life \[(1-r)^{5.5} = \frac{1}{2}\] \[r = 1 - (\frac{1}{2})^{5.5}\]

OpenStudy (dumbcow):

oops just realized i had a typo \[r = 1 - (\frac{1}{2})^{1/5.5}\]

OpenStudy (anonymous):

well the answer is supposed to be .126

OpenStudy (dumbcow):

well in that case it must be an exponential decay \[e^{-5.5k} = \frac{1}{2}\] \[k = \frac{\ln (1/2)}{-5.5}\]